# More Likely Than You Think: Inclination-Driving Secular Resonances are Common in Known Exoplanet Systems

### v1

@@ 5 Discussion
The impact of sweeping resonances via $J_{2}$ also poses an interesting possibility. The size of the sweeping region in semimajor axis ratio space (see Figure 6) depends on the maximum value of $J_{2}$ that the star experiences. This means that the present-day mutual inclinations between planets depend on the maximum $J_{2}$ achieved, meaning that the maximum spin of the star (and therefore the maximum $J_{2}$) could be constrained by present-day planetary configurations. Because stellar spins as functions of time tend to be relatively indistinguishable after $\sim 1\,$Gyr regardless of initial spin (see, e.g., Barnes, 2003; Hartman et al., 2010; Delorme et al., 2011; Meibom et al., 2015), this dynamical channel could provide a way to probe the initial conditions of billions of years-old stars. Simultaneously, the trends in Figure 7 show a remarkable dependence between mean mutual inclination at late times and initial mutual inclination. This relationship may allow some constraints to also be put on the maximum mutual inclination between outer planets during the early years of the system’s life–when $J_{2}$ was moving through its high values that drive the resonances. A straightforward implementation of this is present in Figure 7, represented by the black dotted line: blue dots above this line should not present as co-transiting systems and are thus forbidden.
@@ 6 Conclusion
To determine how common sweeping resonances from evolving $J_{2}$ are, we explored the population of observed, three-planet systems. We found that $\sim 20\%$ of a sample of three-planet all-transiting systems have mass and semimajor axis ratios that indicate they underwent sweeping resonances early in their lives. In Figure 7, we show that the ability of sweeping resonances to misalign the inner planet from its companions is dependent on the initial mutual inclination between the outer two planets. Moreover, for all six sweeping-resonant systems we identified, they would require very large outer planet mutual inclinations to make the inner, transiting planet no longer transit–consistent with observations. Our findings have significant implications for the connection between exoplanet systems’ initial conditions and their present configurations. A fraction of systems experiencing sweeping resonances of near $20\%$ constitutes a generic, ubiquitous phenomenon.

### v2

@@ 5 Discussion
The impact of sweeping resonances via $J_{2}$ also poses an interesting possibility. The size of the sweeping region in semimajor axis ratio space (see Figure 6) depends on the maximum value of $J_{2}$ that the star experiences. This means that the present-day mutual inclinations between planets depend on the maximum $J_{2}$ achieved, meaning that the maximum spin of the star (and therefore the maximum $J_{2}$) could be constrained by present-day planetary configurations. Because stellar spins as functions of time tend to be relatively indistinguishable after $\sim 1\,$Gyr regardless of initial spin (see, e.g., Barnes, 2003; Hartman et al., 2010; Delorme et al., 2011; Meibom et al., 2015), this dynamical channel could provide a way to probe the initial conditions of billions of years-old stars. Simultaneously, the trends in Figure 7 show a remarkable dependence between mean mutual inclination at late times and initial mutual inclination. This relationship may allow some constraints to also be put on the maximum mutual inclination between outer planets during the early years of the system’s life–when $J_{2}$ was moving through its high values that drive the resonances. A straightforward implementation of this is present in Figure 7, represented by the blue dotted line: blue dots above this line should not present as co-transiting systems and are thus forbidden.
@@ 6 Conclusion
To determine how common sweeping resonances from evolving $J_{2}$ are, we explored the population of observed, three-planet systems. We found that $\sim 20\%$ of a sample of three-planet all-transiting systems have mass and semimajor axis ratios that indicate they underwent sweeping resonances early in their lives. In Figure 7, we show that the ability of sweeping resonances to misalign the inner planet from its companions is dependent on the initial mutual inclination between the outer two planets. Moreover, for all six sweeping-resonant systems we identified, they would require very large initial outer planet mutual inclinations to make the inner, transiting planet no longer transit. As such high initial mutual inclinations are unlikely, that these systems are observed in co-transiting configurations is unsurprising. Our findings have significant implications for the connection between exoplanet systems’ initial conditions and their present configurations. A fraction of systems experiencing sweeping resonances of near $20\%$ constitutes a generic, ubiquitous phenomenon.